OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for ballet classes. Let f(x) represent the total number of students who enrolled for the classes after x years. f(x) = 12(1.06)^x The average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is ________students per year. Round your answer to the nearest whole number. I got 0.8832... Do I round up or down (because it's a person)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

How did you get 0.8832?

OpenStudy (anonymous):

I typed 12(1.06)^5-12(1.06)^3/5-3 in my calculator

OpenStudy (anonymous):

Actually I got 1.77

OpenStudy (anonymous):

12*.1.06^3 = 14.29 12*1.06^5 = 16.06 16.06-14.29 = 1.77

OpenStudy (anonymous):

Nearest whole number would be 2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

did that help you out?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok :)